# TAROT: Targeted Data Selection via Optimal Transport

### v1

@@ Abstract
We propose TAROT, a Targeted data selection framework grounded in Optimal Transport theory. Previous targeted data selection methods primarily use influence-based greedy heuristics to enhance domain-specific performance. These methods perform well on limited, unimodal data (i.e., data following a single pattern) but become less effective as target data increases in complexity. Specifically, in multimodal distributions, these heuristics fail to account for multiple inherent patterns, leading to suboptimal data selection. This work identifies two primary factors contributing to this limitation: (i) the disproportionate impact of dominant feature components in high-dimensional influence estimation, and (ii) the restrictive linear additive assumptions inherent in greedy selection strategies. To address these challenges, TAROT incorporates whitened feature distance to mitigate dominant feature bias, offering a more reliable measure of data influence. Building on this, TAROT uses whitened feature distance to quantify and minimize the optimal transport distance between the selected data and target domains. Notably, this minimization also facilitates the estimation of optimal selection ratios. We evaluate TAROT across multiple tasks, including semantic segmentation, motion prediction, and instruction tuning. Results consistently show that TAROT outperforms state-of-the-art methods, highlighting its versatility across various deep learning tasks. Code is available at https://github.com/vita-epfl/TAROT.
Figure 1: Left: this figure demonstrates the selection objective of TAROT, designed to minimize optimal transport costs. The method operates on a target distribution and a candidate data pool, selecting a subset from the candidate data to reduce transport costs. Right: the figure highlights the approximate performance improvements and data efficiency achieved by TAROT across various tasks. It illustrates that TAROT consistently enhances task performance in different domains while requiring less data. Detailed experimental results are provided in Sec. 5.
@@ 1 Introduction
Expanding model capacity and data volume has emerged as a key strategy in both computer vision and large language models [1, 47, 33, 15]. However, training a model on all available data may not be optimal or feasible, as data quality varies and computational resources are often limited [2, 29, 13, 25]. To address this, recent research has introduced targeted data selection methods. These methods aim to select data from a candidate pool that maximizes model performance on specific target tasks. Recent methods [42, 13] address this challenge using a two-step approach: (1) influence estimation, which calculates data influence scores between the candidate and target datasets, and (2) greedy heuristics, which selects data with the highest average influence scores for the target dataset.
@@
Although these selection paradigms have shown to be effective in large language model (LLM) applications [13, 42], they rely heavily on the assumption that influence functions are additive. This means that the influence of multiple training samples on a target sample can be added together. However, as noted by Hu et al. [22], this assumption, while efficient and broadly applicable, fails even in simple linear regression scenarios. One main reason why the additivity assumption works well in some cases is the simple and unimodal distribution of the target domain, such as question-answer pairs from a specific domain in LLMs. In these situations, adding up the influences provides a good estimate of how a candidate sample impacts the entire target dataset. However, in other contexts like motion prediction, where data distributions are often complex and multimodal, this approach falls short. In such cases, the summed influence does not capture the diverse nature of the target distribution, reducing its effectiveness.
@@
To this end, we adopt a distribution-matching approach for targeted data selection. We first introduce whitened feature distance (WFD), an influence estimation method designed to mitigate the bias from dominant features. Using WFD as pairwise cost estimation between data points, we develop TAROT, an Optimal Transport (OT)-based targeted selection framework. This method selects data that minimizes the OT distance between the target distribution and the selected samples. Fig. 1 illustrates the selection objective of our approach. Our choice of OT is also supported by findings in [24], which demonstrate that training models on source data with smaller OT distances to the target domain improve domain-specific performance.
@@
To validate the general effectiveness of our method, we conduct experiments on semantic segmentation, motion prediction, and instruction tuning tasks. Extensive results demonstrate that our approach consistently outperforms previous methods. Notably, models trained on smaller, domain-relevant datasets often surpass those trained on larger, less relevant datasets.
In summary, our contributions are as follows:
@@
  • Broad applicability and improved performance: extensive experiments in computer vision and LLMs demonstrate that TAROT consistently outperforms existing methods, allowing models trained on smaller, targeted datasets to match or surpass the performance of those trained on full datasets.
2 Related Work
Data Influence Estimation. Pruthi et al. [35] proposed a method to estimate data influence by tracing the gradient descent process. Another approach utilizes datamodels: Ilyas et al. [23] showed that simple linear datamodels can predict model outputs effectively. Building on this, Park et al. [34] introduced a scalable technique for estimating datamodel scores. Subsequent works have optimized these methods, focusing on checkpoint ensembles [11], efficient gradient storage [46], and scalable gradient projection [7]. Our method extends Pruthi et al. [35] by enhancing estimation robustness through feature covariance whitening.
Coresets and Targeted Selection. Coreset selection aims to identify a dataset subset for optimal in-domain performance [43, 31, 4, 27, 17, 44], assuming validation and candidate data share the same distribution. Our work shifts focus to targeted selection, optimizing performance on a distinct target domain. Prior methods often rely on influence-estimation heuristics: Engstrom et al. [13] leveraged linear datamodel scores [34] to select data for LLMs pretraining; Xia et al. [42] leveraged influence estimation to select data for LLMs fine tuning. Hu et al. [22] highlighted these methods’ limitations and proposed iterative heuristics. Our approach introduces a novel distribution-matching framework to address these challenges.
Optimal Transport for Data Selection. Optimal transport has been used for dataset similarity estimation [3] and sensitivity analysis in data valuation [24]. Kang et al. [25] applied it to optimize data selection ratios across multiple sources. We integrate influence estimation with optimal transport, enabling effective subset selection to boost target domain performance.
@@ 3.1 Data Influence Formulation
Following Pruthi et al. [35], we summarize how to estimate the influence of a training data point on held-out data using a first-order approximation of training dynamics.
First-order Approximation to Per-step Influence. Suppose $x_{i}$ is the only training sample at time step $t$, and the model $\theta^{t}$ is updated based on the loss $\mathcal{L}(x_{i};\theta^{t})$ using SGD with learning rate $\eta_{t}$. The per-step influence of $x_{i}$ on a target sample $x$ is defined as the reduction in the loss of $x$ caused by training on $x_{i}$. This can be approximated using a first-order Taylor expansion. The per-step influence is given by:
@@
Trajectory Influence via Checkpoints. For practicality, Pruthi et al. [35] compute the trajectory influence by summing the per-step influences over epochs:
@@ 3.2 Background on Optimal Transport
Given a complete, separable metric space $\mathcal{X}$ and probability measures $\alpha,\beta\in\mathcal{P}(\mathcal{X})$, the Kantorovich formulation [26] of OT is defined as:
@@
In practice, the measures $\alpha$ and $\beta$ are approximated using finite samples $\{\mathbf{x}^{(i)}\}$ and $\{\mathbf{y}^{(j)}\}$, with corresponding discrete measures $\alpha=\sum_{i=1}^{n}a_{i}\delta_{\mathbf{x}^{(i)}}$ and $\beta=\sum_{j=1}^{m}b_{j}\delta_{\mathbf{y}^{(j)}}$. The pairwise costs form an $n\times m$ matrix $\mathbf{C}$, where $\mathbf{C}_{ij}=c(\mathbf{x}^{(i)},\mathbf{y}^{(j)})$. Solving this problem as a linear program is computationally expensive, with cubic complexity. In implementation, we use the Sinkhorn algorithm [10] for efficient calculation.
@@ 4 Method
The failure modes of influence-based greedy heuristics have been systematically studied in [22]. Specifically, errors in the influence function and the inability to account for the non-additive nature of collective influence can cause these heuristics to fail even in simple linear regression tasks.
In contrast, we frame targeted data selection as a distribution matching problem, aiming to select a subset of candidate data that closely aligns with the target distribution. We use the OT distance to quantify the discrepancy between distributions.
Secs. 4.1 and 4.2 discuss how to measure the OT distance between datasets, and Sec. 4.3 discusses how to select data to minimize the OT distance. Fig. 2 shows an overview of these two processes.
@@ 4.1 Optimal Transport Between Datasets
As discussed in [3], estimating OT distance between datasets begins with defining a distance metric between data pairs $(x,y)$ and $(x^{\prime},y^{\prime})$. In Sec. 3.1, we introduced a gradient-based method for estimating data influence, which embeds both input features and labels into the model’s gradient. This gradient feature captures the task-specific relationship between data points and the model’s loss. By computing distances such as the L2 distance, it provides a viable metric for the OT problem. Furthermore, it aligns with the model’s optimization objective, ensuring that the OT distance reflects meaningful differences between datasets in terms of their impact on model training.
@@ 4.2 Whitened Feature Distance
Directly computing the distance $d_{\mathcal{Z}}\left(z,z^{\prime}\right)$ as defined in Eq. 5 presents significant challenges due to the correlation of gradient features. This correlation results in an ill-conditioned covariance matrix, allowing certain feature space directions to dominate distance estimation and distort sample relationships. Additionally, varying gradient component scales bias distance measurements, impairing the accuracy of distance estimation.
@@
    Random projection. We adopt the same random projection technique from [42, 34] to reduce the dimensionality of the gradients. The projected gradient $\phi(z)^{\text{proj}}_{i}$ is computed as:
@@
    where $\mathcal{P}^{T}$ is the projection matrix. We leverage the efficient CUDA implementation from Park et al. [34].
@@
    Next, we compute the covariance matrix:
    $\Sigma=\frac{1}{N}\sum_{i=1}^{N}\tilde{\phi}(z)_{i}^{\text{proj}}(\tilde{\phi}(z)_{i}^{\text{proj}})^{\top},$ (8)
    and apply Cholesky decomposition, $\Sigma=LL^{\top}$, where $L$ is lower triangular. The gradients are whitened as:
@@
This gradient-based distance also applies to data influence estimation. Our experiments show that WFD outperforms the state-of-the-art method [34], as detailed in Sec. 5.1.
@@ 4.3 Data Selection via OT Distance Minimization
where $\Pi(\mathcal{D}_{s}\cup\{z_{i}\},\mathcal{D}_{t})$ is the set of couplings between $\mathcal{D}_{s}\cup\{z_{i}\}$ and $\mathcal{D}_{t}$. As noted in [24], $\phi_{i}$ estimates the benefit of adding $z_{i}$ for minimizing the OT distance. We rank candidates by potential and select the top $S-|\mathcal{D}_{s}|$ samples. The process ends once these samples are added.
@@ 4.4 Data Weighting with OT Potential
To highlight the importance of individual data points, we assign weights based on their OT potentials from Eq. 14. These potentials are scaled to positive integer weights, determining the number of times each point is duplicated within an epoch. The scaled weights $\{w_{i}\}$ are normalized to satisfy $\sum_{i=1}^{S}w_{i}=R$, where $R$ is a customizable repetition factor. For instance, $R$ can be set to $N+M$ to match the same training steps as using the full dataset, or adjusted based on specific training requirements, e.g., limited training budget.
@@ 5 Experiments
This section starts by evaluating the influence estimation performance of WFD in Sec. 5.1. Following that, we evaluate TAROT across diverse tasks, including semantic segmentation (Sec. 5.2), motion prediction (Sec. 5.3), and instruction tuning (Sec. 5.4).
@@ 5.1 WFD for Influence Estimation
Evaluation Metric. Following Park et al. [34], we use the Linear Data Modeling Score (LDS) as our evaluation metric. LDS measures the effectiveness of an attribution or influence estimation method in predicting changes in model outputs when the training set is altered. Further details of LDS are provided in the supplementary materials.
Target and Candidate Datasets. Following Park et al. [34], we evaluate image classification using ResNet-9 classifiers trained on the CIFAR-10 dataset. For motion prediction, we adopt AutoBots [16], training on the nuScenes [5] dataset (32k samples) and validating on 9k target samples.
Baselines. We compare WFD against the state-of-the-art data attribution method TRAK [34] using the same ensemble approach. Since WFD quantifies distance (where smaller values indicate greater influence), we evaluate it with the LDS score by considering negative WFD values. Additionally, we employ ZCA whitening instead of Cholesky whitening to preserve data structure for improved LDS performance. Differences between these two whitening methods are discussed in the supplementary materials.
Figure 4: WFD outperforms baseline methods on both tasks. The x-axis indicates the number of model checkpoints that a given method uses to compute attribution scores. The y-axis indicates the method’s efficacy as measured by the linear datamodeling score (LDS).
Main Results. Fig. 4 compares the performance of WFD with TRAK on the LDS score. Key findings include:
@@
  • Inspecting WFD-identified examples. In  Fig. 3 we display, for four randomly selected target samples from the nuScenes [5] and Cityscapes [9] dataset, the training examples from WOMD [14] and GTA5 [36] dataset corresponding to the smallest (most positive) and largest (most negative) whitened feature distance.
@@ 5.2 TAROT for Semantic Segmentation
Semantic segmentation, a key task in computer vision, involves assigning semantic labels to every pixel in an image. Prior studies show that synthetic data can significantly boost segmentation model performance on real-world data [36]. In this section, we address two research questions:
  • Is all synthetic data equally effective for improving real-world segmentation performance?
  • Can targeted selection methods identify synthetic data that better align with real-world distributions to enhance model performance?
Target and Candidate Datasets. The GTA5 dataset [36] serves as the candidate dataset, while the Cityscapes [9] training split (2975 samples) is used as the target dataset, with its validation split for evaluation. GTA5 contains 24,966 synthetic images with pixel-level annotations of urban scenes, whereas Cityscapes provides finely annotated real-world urban images.
@@
Figure 6: Data selection results visualized in the gradient feature space using T-SNE. We use Cityscapes’ validation set as the target, we select 10% of its training set. Left: DsDm, where selected samples cluster tightly around the target distribution center. Right: TAROT, where selected samples are distributed more broadly, capturing the complexity of the target distribution.
Baselines. We compare TAROT with state-of-the-art targeted selection methods LESS and DsDm [42, 13], as well as random selection.
Data Selection and Model Training. We employ DeepLabV3 [6] with a ResNet50 [19] backbone, using gradients from the last four checkpoints to guide data selection. We evaluate selection ratios of 5%, 20%, 50% and OTM (Sec. 4.3). OTM selects approximately 24% of the data. The repetition factor is set as: $R=N+M$ for data selected by TAROT (Sec. 4.4), while other methods employ uniform repetition to match the same size.
A new DeepLabV3 model is then trained from scratch on the selected subsets for 50 epochs with a batch size of 16. No data augmentation is applied, and images are resized to $920\times 460$.
@@
Main Results. Fig. 5 compares the performance of TAROT and other baselines. Key findings include:
@@
  • Additional insights from T-SNE visualizations. Fig. 6 shows T-SNE visualizations of gradient features for selected data. The visualizations reveal differing selection behaviors: DsDm tightly clusters around the target distribution center, failing to capture its complexity. In contrast, TAROT selects data that better aligns with the broader target distribution, highlighting a key limitation of averaging-based influence methods in handling complex distributions.

### v2

@@ Abstract
We propose TAROT, a Targeted data selection framework grounded in Optimal Transport theory. Previous targeted data selection methods primarily use influence-based greedy heuristics to enhance domain-specific performance. These methods perform well on limited, unimodal data (i.e., data following a single pattern) but become less effective as target data increases in complexity. Specifically, in multimodal distributions, these heuristics fail to account for multiple inherent patterns, leading to suboptimal data selection. This work identifies two primary factors contributing to this limitation: (i) the disproportionate impact of dominant feature components in high-dimensional influence estimation, and (ii) the restrictive linear additive assumptions inherent in greedy selection strategies. To address these challenges, TAROT incorporates whitened feature distance to mitigate dominant feature bias, offering a more reliable measure of data influence. Building on this, TAROT uses whitened feature distance to quantify and minimize the optimal transport distance between the selected data and target domains. Notably, this minimization also facilitates the estimation of optimal selection ratios. We evaluate TAROT across multiple tasks, including semantic segmentation, motion prediction, and instruction tuning. Results consistently show that TAROT outperforms state-of-the-art methods, highlighting its versatility across various deep learning tasks. Code is available at: https://github.com/vita-epfl/TAROT.
Data Attribution, Data Selection, Optimal Transport
Figure 1: Left: the selection objective of TAROT, designed to minimize optimal transport costs. The method operates on a target distribution and a candidate data pool, selecting a subset from the candidate data to reduce transport costs. Right: the approximate performance improvements and data efficiency achieved by TAROT across various tasks. It illustrates that TAROT consistently enhances task performance in different domains while requiring less data. Detailed experimental results are provided in Section 4.
@@ 1 Introduction
Expanding model capacity and data volume has emerged as a key strategy in both computer vision (CV) and large language models (LLMs) (Achiam et al., 2023; Zou et al., 2024; Oquab et al., 2023; Feng et al., 2024). However, training a model on all available data may not be optimal or feasible, as data quality varies and computational resources are often limited (Albalak et al., 2024; Liu et al., 2024; Engstrom et al., 2024; Kang et al., 2024). To address this, recent research has introduced targeted data selection methods. These methods aim to select data from a candidate pool that maximizes model performance on specific target tasks. Recent methods (Xia et al., ; Engstrom et al., 2024) address this challenge using a two-step approach: (1) influence estimation, which calculates data influence scores between the candidate and target datasets, and (2) greedy heuristics, which selects data with the highest average influence scores for the target dataset.
@@
Although these selection paradigms have shown to be effective in LLM applications (Engstrom et al., 2024; Xia et al., ), they rely heavily on the assumption that influence functions are additive. This means that the influence of multiple training samples on a target sample can be added together. However, as noted by Hu et al. (2024), this assumption, while efficient and broadly applicable, fails even in simple linear regression scenarios. One main reason why the additivity assumption works well in some cases is the simple and unimodal distribution of the target domain, such as a single task (e.g. code generation) in LLMs. In these situations, adding up the influences provides a good estimate of how a candidate sample impacts the entire target task. However, in other contexts like motion prediction, where data distributions are often complex and multimodal, this approach falls short. In such cases, the summed influence does not capture the diverse nature of the target distribution, reducing its effectiveness.
@@
To this end, we adopt a distribution-matching approach for targeted data selection. We first introduce whitened feature distance (WFD), an influence estimation method designed to mitigate the bias from dominant features. The benefits of feature whitening are also discussed in (Ermolov et al., 2021). Using WFD as pairwise cost estimation between data points, we develop TAROT, an Optimal Transport (OT)-based targeted selection framework. This method selects data that minimizes the OT distance between the target distribution and the selected samples. Figure 1 illustrates the selection objective of our approach. Our choice of OT is also supported by findings in (Just et al., 2023), which demonstrate that training models on source data with smaller OT distances to the target domain improves domain-specific performance.
@@
To validate the general effectiveness of our method, we conduct experiments on semantic segmentation, motion prediction, and instruction tuning tasks. Extensive results demonstrate that our approach consistently outperforms previous methods. Notably, models trained on smaller, domain-relevant datasets often surpass those trained on larger, less relevant datasets. In summary, our contributions are:
@@
  • Broad applicability and improved performance: extensive experiments in CV and LLMs demonstrate that TAROT consistently outperforms existing methods, allowing models trained on smaller, targeted datasets to match or surpass the performance of those trained on full datasets.
@@ 2.1 Data Influence Formulation
Following Pruthi et al. (2020), we summarize how to estimate the influence of a training data point on held-out data using a first-order approximation of training dynamics.
First-order Approximation to Per-step Influence. Suppose $x_{i}$ is the only training sample at time step $t$, and the model $\theta^{t}$ is updated based on the loss $\mathcal{L}(x_{i};\theta^{t})$ using SGD with learning rate $\eta_{t}$. The per-step influence of $x_{i}$ on a target sample $x$ is defined as the reduction in the loss of $x$ caused by training on $x_{i}$. This can be approximated using a first-order Taylor expansion. The per-step influence is:
@@
Trajectory Influence via Checkpoints. For practicality, Pruthi et al. (2020) compute the trajectory influence by summing the per-step influences over epochs:
@@ 2.2 Background on Optimal Transport
Given a complete, separable metric space $\mathcal{X}$ and probability measures $\alpha,\beta\in\mathcal{P}(\mathcal{X})$, the Kantorovich formulation (Kantorovitch, 1958) of OT is defined as:
@@
In practice, the measures $\alpha$ and $\beta$ are approximated using finite samples $\{\mathbf{x}^{(i)}\}$ and $\{\mathbf{y}^{(j)}\}$, with corresponding discrete measures $\alpha=\sum_{i=1}^{n}a_{i}\delta_{\mathbf{x}^{(i)}}$ and $\beta=\sum_{j=1}^{m}b_{j}\delta_{\mathbf{y}^{(j)}}$. The pairwise costs form an $n\times m$ matrix $\mathbf{C}$, where $\mathbf{C}_{ij}=c(\mathbf{x}^{(i)},\mathbf{y}^{(j)})$. Solving this problem as a linear program is computationally expensive, with cubic complexity. In implementation, we use the Sinkhorn algorithm (Cuturi, 2013) for efficient calculation.
@@ 3 Method
The failure modes of influence-based greedy heuristics have been systematically studied in (Hu et al., 2024). Specifically, errors in the influence function and the inability to account for the non-additive nature of collective influence can cause these heuristics to fail even in simple linear regression tasks.
In contrast, we formulate targeted data selection as a distribution matching problem, aiming to select a subset of candidate data that closely aligns with the target distribution. We employ the OT distance to quantify the discrepancy between distributions.
Sections 3.1 and 3.2 demonstrate how to measure the OT distance between datasets, and Section 3.3 discusses strategies for selecting data to minimize the OT distance. An overview of these processes is illustrated in Figure 2.
@@ 3.1 Optimal Transport Between Datasets
As discussed in (Alvarez-Melis & Fusi, 2020), estimating OT distance between datasets begins with defining a distance metric between data pairs $(x,y)$ and $(x^{\prime},y^{\prime})$. In Section 2.1, we introduced a gradient-based method for estimating data influence, which embeds both input features and labels into the model’s gradient. This gradient feature captures the task-specific relationship between data points and the model’s loss. By computing distances such as the L2 distance, it provides a viable metric for the OT problem. Furthermore, it aligns with the model’s optimization objective, ensuring that the OT distance reflects meaningful differences between datasets in terms of their impact on model training.
@@ 3.2 Whitened Feature Distance
Directly computing the distance $d_{\mathcal{Z}}\left(z,z^{\prime}\right)$ as defined in Equation 5 presents significant challenges due to the correlation of gradient features. This correlation results in an ill-conditioned covariance matrix, allowing certain feature space directions to dominate distance estimation and distort sample relationships. Additionally, varying gradient component scales bias distance measurements, impairing the accuracy of distance estimation.
@@
    Random projection. We adopt the same random projection technique from (Xia et al., ; Park et al., 2023) to reduce the dimensionality of the gradients. The projected gradient $\phi(z)^{\text{proj}}_{i}$ is computed as:
@@
    where $\mathcal{P}^{T}$ is the projection matrix. We leverage the efficient CUDA implementation from Park et al. (2023).
@@
    Next, we compute the covariance matrix $\Sigma$ and apply Cholesky decomposition, $\Sigma=LL^{\top}$, where $L$ is lower triangular. The gradients are whitened as:
@@
This gradient-based distance also applies to data influence estimation. Our experiments show that WFD outperforms the state-of-the-art method (Park et al., 2023), as detailed in Section 4.1.
@@ 3.3 Data Selection via OT Distance Minimization
where $\Pi(\mathcal{D}_{s}\cup\{z_{i}\},\mathcal{D}_{t})$ is the set of couplings between $\mathcal{D}_{s}\cup\{z_{i}\}$ and $\mathcal{D}_{t}$. As noted in (Just et al., 2023), $\phi_{i}$ estimates the benefit of adding $z_{i}$ for minimizing the OT distance. We rank candidates by potential and select the top $S-|\mathcal{D}_{s}|$ samples. The process ends once these samples are added.
@@ 3.4 Data Weighting with OT Potential
To highlight the importance of individual data points, we assign weights based on their OT potentials from Equation 13. These potentials are scaled to positive integer weights, determining the number of times each point is duplicated within an epoch. The scaled weights $\{w_{i}\}$ are normalized to satisfy $\sum_{i=1}^{S}w_{i}=R$, where $R$ is a customizable repetition factor. For instance, $R$ can be set to $N+M$ to match the same training steps as using the full dataset, or adjusted based on specific training requirements, e.g., limited training budget.
@@ 4 Experiments
This section starts by evaluating the influence estimation performance of WFD in Section 4.1. Following that, we evaluate TAROT across diverse tasks, including semantic segmentation (Section 4.2), motion prediction (Section 4.3), and instruction tuning (Section 4.4).
@@ 4.1 WFD for Influence Estimation
Evaluation Metric. Following Park et al. (2023), we use the Linear Data Modeling Score (LDS) as our evaluation metric. LDS measures the effectiveness of an attribution or influence estimation method in predicting changes in model outputs when the training set is altered. Further details of LDS are provided in the supplementary materials.
Target and Candidate Datasets. Following Park et al. (2023), we evaluate image classification using ResNet-9 classifiers trained on the CIFAR-10 dataset. For motion prediction, we adopt AutoBots (Girgis et al., 2021), training on the nuScenes (Caesar et al., 2020) dataset (32k samples) and validating on 9k target samples.
Baselines. We compare WFD against the state-of-the-art data attribution method TRAK (Park et al., 2023) using the same ensemble approach. Since WFD quantifies distance (where smaller values indicate greater influence), we evaluate it with the LDS score by considering negative WFD values. Additionally, we employ ZCA whitening instead of Cholesky whitening to preserve data structure for improved LDS performance. Differences between these two whitening methods are discussed in the supplementary materials.
Figure 4: WFD outperforms baseline methods on both tasks. The x-axis indicates the number of model checkpoints that a given method uses to compute attribution scores. The y-axis indicates the method’s efficacy as measured by LDS.
Main Results. Figure 4 compares the performance of WFD with TRAK on the LDS score. Key findings include:
@@
  • Inspecting WFD-identified examples. In  Figure 3 we display, for four randomly selected target samples from the nuScenes (Caesar et al., 2020) and Cityscapes (Cordts et al., 2016) dataset, the training examples from WOMD (Ettinger et al., 2021) and GTA5 (Richter et al., 2016) dataset corresponding to the smallest (most positive) and largest (most negative) whitened feature distance.
@@ 4.2 TAROT for Semantic Segmentation
Semantic segmentation involves assigning semantic labels to every pixel in an image. Prior studies show that synthetic data can significantly boost segmentation model performance on real-world data (Richter et al., 2016). We then address the following research question: Can TAROT identify synthetic data that better align with real-world distributions to enhance model performance?
Datasets. The GTA5 dataset (Richter et al., 2016) serves as the candidate dataset, while the Cityscapes (Cordts et al., 2016) training split (2975 samples) is used as the target dataset, with its validation split for evaluation. GTA5 contains 24,966 synthetic images with pixel-level annotations of urban scenes, whereas Cityscapes provides finely annotated real-world urban images.
@@
Figure 6: Data selection visualization in gradient feature space with T-SNE. We select 10% of Cityscapes’ training set, with validation split as the target. Left: Samples selected by DsDm cluster tightly around the target distribution center. Right: Samples selected by TAROT are distributed more broadly, capturing the target distribution’s complexity.
Baselines. We compare TAROT with state-of-the-art targeted selection methods LESS and DsDm (Xia et al., ; Engstrom et al., 2024), as well as random selection.
Data Selection and Model Training. We employ DeepLabV3 (Chen, 2017) with a ResNet50 (He et al., 2016) backbone, using gradients from the last four checkpoints to guide data selection. We evaluate selection ratios of 5%, 20%, 50% and OTM (Section 3.3). OTM selects approximately 24% of the data. The repetition factor is set as: $R=N+M$ for data selected by TAROT (Section 3.4), while other methods employ uniform repetition to match the same size. A new DeepLabV3 model is then trained from scratch on the selected subsets.
@@
Main Results. Figure 5 compares the performance of TAROT and other baselines. Key findings include:
@@
  • Additional insights from T-SNE visualizations. Figure 6 shows T-SNE visualizations of gradient features for selected data. The visualizations reveal differing selection behaviors: DsDm tightly clusters around the target distribution center, failing to capture its complexity. In contrast, TAROT selects data that better aligns with the broader target distribution, highlighting a key limitation of averaging-based influence methods in handling complex distributions.